Matt is 4 years older than Todd. In 3 years. Twice matt’s Age will be 8 more than twice Todd’s age. Find their ages

Answer by ikleyn(52824)

.

Consider punctuation in your post.

Look in it so long until you find an error in it.

Then fix your error.

Then re-post.


Have a nice day . . .


----------------
Comment from student: Matt is 4 years older than Todd. In 3 years , twice matt’s Age will be 8 more than twice Todd’s age. Find their ages.
----------------


My response : OK.

    M = T + 4                (1)     ("Matt is 4 years older than Todd.")

    2*(M+3) = 2*(T+3) + 8    (2)     ("In 3 years , twice Matt's age will be 8 more than twice Todd’s age.")


First step simplify (2)


    2M + 6 = 2T + 6 + 8

    2M = 2T + 8

    M =  T + 4                (3)


Thus you have THIS:


    equation (2) is equivalent to equation (3), which is identical to equation (1).


It means that equations (1) and (3) are identical and have infinitely many solutions, do not allowing to select one unique solution.


It means, in turn, that the problem condition is  DEFECTIVE.


It also means that I spent my time for nothing working on your defective condition.